If a baseball player throws a ball at 35.0m/s, what is its maximum range?
1
Expert's answer
2019-11-11T15:52:38-0500

The maximum range:


"R=\frac{v^2}{g}"

"R=\frac{35^2}{9.8}=125\ m"